OpenStudy (anonymous):

What is the graph of the function rule? y=I4xI+3 Which way will the lines point up or down

OpenStudy (anonymous):

\[y-3=\left| 4x \right|\] \[y-3=4x ,if x \ge 0 ,i.e.,y-3\ge 0 or y \ge 3\] \[y-3=-4x ,if x <0 or R.H.S. is positive.\] \[y-3 \ge 0 or y \ge 3\] \[In either case y \ge 0\] so you see lines are upward.

OpenStudy (anonymous):

Here's a way of thinking about it. The absolute value of 4x tells you no matter what x value you put in, positive or negative, |4x| will always be greater than 0. Obviously the +3 is greater than 0 as well. So the graph is pointing upwards.

OpenStudy (anonymous):

would the lines go up from 3 or -3 @surjithayer

OpenStudy (anonymous):

from3

OpenStudy (anonymous):

How the graph would look. y=4x would just be a straight line going from bottom left to top right. y=|4x| would be the same, except when the x values become negative you get positive values mirroring the other side. So the line would bend at the origin and you would get two lines that reflect each other. y=x+3 has the effect of shifting the entire graph of y=x up by 3 points. So in y=|4x|+3 it would shift the graph of y=|4x| up by 3 points.

OpenStudy (ranga):